OpenStudy (anonymous):

2.4 written as a mixed number in simplest form

OpenStudy (akashdeepdeb):

2.4 written as a fraction is 24/10 So how many times does 10 go in 24? 2 times So it has to be 2 whole and 4/10 2 4/10 SIMPLIFY THAT! :D 2 2/5 Understood? :)

OpenStudy (anonymous):

\[2\frac{ 2 }{ 5 }\] and \[\frac{ 12 }{ 5 }\] Write the decimal number as a fraction (over 1) 2.4 = 2.4 / 1 Multiplying by 1 to eliminate 1 decimal places we multiply top and bottom by 10. For ever number behind the decimal you'd add a zero to 10. In this case there's only one behind it so we multiply by 10. :) Numerator (N) N = 2.4 × 10 = 24 Denominator (D) D = 1 × 10 = 10 N / D = 24 / 10 \[\frac{ 2.4 }{ 1}\times \frac{ 10 }{ 10}=\frac{ 24 }{ 10 }=\frac{ 12 }{ 5}\] Hope I was of great help. Cheers!